# Cohomological Calibration and Curvature Constraints on Product Manifolds: A Topological Lower Bound

### v4

@@ 5. Main Result and Proof
$\Psi_{p}:\mathrm{span}\{\alpha_{i}\otimes\beta_{i}\}\longrightarrow\mathfrak{so}(T_{p}M)$
whose injectivity follows from the pointwise independence of the evaluations $\{\alpha_{i}(p_{1})\}$ and $\{\beta_{i}(p_{2})\}$: any nontrivial relation among the images would imply a corresponding relation among the tensor products $\alpha_{i}(p_{1})\otimes\beta_{i}(p_{2})$, forcing all coefficients to vanish. Hence $\Psi_{p}$ produces $r$ linearly independent endomorphisms $R^{\mathrm{off}}(X_{i},Y_{i})Z_{i}$ in $\mathfrak{so}(T_{p}M)$.
The linear map
$\Psi_{p}:\mathrm{span}_{\mathbb{R}}\{\alpha_{i}\otimes\beta_{i}\}_{i=1}^{r}\to\mathfrak{so}(T_{p}M)$
defined by
$\Psi_{p}(\alpha_{i}\otimes\beta_{i})(Z):=\tfrac{1}{2}(\nabla^{LC}T)^{\mathrm{off}}(X,Y,Z)$
[…]
The map $\Psi_{p}$ is defined by the linear part of the curvature, which is proportional to $(\nabla^{LC}T)_{\text{off}}$. The condition that a linear combination of these contributions vanishes at a generic point $p=(p_{1},p_{2})$ is:
$\sum_{i}c_{i}\left[(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}+\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i})\right]_{p}=0$
Since the forms $\alpha_{i}$ and $\beta_{i}$ are linearly independent and the covariant derivatives $\nabla^{g_{1}}$ and $\nabla^{g_{2}}$ act only on their respective factors, the vanishing of the mixed components of this sum implies that the terms must vanish separately.
Specifically, considering the component of type $(2,1)$ (forms with two indices on $M_{1}$ and one on $M_{2}$), the vanishing condition implies:
$\sum_{i}c_{i}(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}=0$
Since the forms $\beta_{i}$ are linearly independent, this requires that the coefficients of $\beta_{i}$ must vanish:
$\sum_{i}c_{i}(\nabla^{g_{1}}\alpha_{i})=0$
This condition holds on an open dense subset $\mathcal{U}\subset M$ (by Lemma 5.1). Since $\alpha_{i}$ are harmonic forms, the unique continuation principle implies that the linear combination $\sum c_{i}\alpha_{i}$ is a $\nabla^{g_{1}}$-parallel 2-form on $M_{1}$. By definition, the space of such forms is $\mathcal{P}_{2}$.
$\sum_{i}c_{i}\alpha_{i}\in\mathcal{P}_{2}$
This proves that the kernel of the map $\Psi_{p}$ is exactly the subspace generated by the simple tensors whose $\alpha$ component is $\nabla^{g_{1}}$-parallel.
Therefore, the kernel of the map $\Psi_{p}$ is precisely the intersection of the subspace spanned by the minimal decomposition with the space of parallel components:
$\ker\Psi_{p}=\left(\text{span}\{\alpha_{i}\otimes\beta_{i}\}_{i=1}^{r_{2,1}}\right)\cap\left(\mathcal{P}_{2}\otimes H^{1}(M_{2})\right)$
Since $\{\alpha_{i}\otimes\beta_{i}\}$ forms a basis for the mixed component $\pi_{2,1}([\omega])$, and the $\alpha_{i}$ are linearly independent, we can choose the basis such that the first $s=\dim(\mathcal{P}_{2})$ elements of $\{\alpha_{i}\}$ are a basis for $\mathcal{P}_{2}$. Given the linear independence of the $\beta_{i}$, the dimension of the kernel is precisely the number of simple tensors in the minimal decomposition whose $\alpha$-component is parallel. This dimension is $\dim(\ker\Psi_{p})=\dim\left(\pi_{2,1}([\omega])\cap(\mathcal{P}_{2}\otimes H^{1}(M_{2}))\right)$.
This establishes the dimension of the off-diagonal curvature generated by the linear term $(\nabla^{\mathrm{LC}}T)_{\mathrm{off}}$ on the open-dense subset $\mathcal{U}^{\prime}$ given by Lemma 5.1.
@@
Then the off-diagonal curvature satisfies $R_{\text{off}}(s)=s(\nabla^{LC}T)_{\text{off}}+s^{2}Q_{\text{off}}$. Differentiating at $s=0$, we see that the image of $(\nabla^{LC}T)_{\text{off}}$ has dimension at least $r^{\sharp}$. The key step is to ensure that the rank of $R_{\text{off}}(s)$ at $s=1$ is not reduced by the quadratic term $Q_{\text{off}}$.
We consider the one-parameter family of connections $\nabla^{(s)}=\nabla^{LC}+s\frac{1}{2}T$. The off-diagonal curvature is given by $R_{\mathrm{off}}(s)=sA_{1}+s^{2}A_{2}$, where $A_{1}=\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ and $A_{2}=Q_{\mathrm{off}}$. On the open-dense set $U$ (defined in Lemma 5.1), the matrix $A_{1}$ has rank $\mathrm{rank}(A_{1})=r^{\sharp}$.
@@
To prove that $\mathrm{rank}(R_{\mathrm{off}}(1))\geq r^{\sharp}$ for the smooth ($C^{\infty}$) metric $g$, we employ the method of analytic approximation and stability of rank (as stated in the Remark 5.6).
By the density theorem of Greene–Jacobowitz [4] and the analyticity of harmonic forms (Morrey–Nirenberg [8]), there exists a sequence of real-analytic metrics $g_{k}\to g$ in the $C^{\infty}$-topology such that $\dim\mathcal{P}_{2}(g_{k})=\dim\mathcal{P}_{2}(g)$ for every $k$; this is possible because the stratum $\{g^{\prime}\mid\dim\mathcal{P}_{2}(g^{\prime})=\dim\mathcal{P}_{2}(g)\}$ is open in the $C^{\infty}$-topology and contains real-analytic metrics densely (Nash–Moser, [9]). Consequently the curvature $R_{\mathrm{off},k}(s)$ associated with $g_{k}$ is real-analytic in $(s,p)$ and the integer $r^{\sharp}$ is the same for every $g_{k}$.
@@
For each fixed $k$, the matrix $R_{\mathrm{off},k}(s)$ is analytic. Since $\mathrm{rank}(A_{1,k})=r^{\sharp}$ on $U$, there exists an $r^{\sharp}\times r^{\sharp}$ minor $M_{k}(s)$ of $R_{\mathrm{off},k}(s)$ such that $M_{k}(s)=s^{r^{\sharp}}D_{k}(s)$, where $D_{k}(0)\neq 0$. By the Identity Principle for Analytic Functions, $D_{k}(s)$ is not identically zero, and its set of zeros is discrete. Thus, $\mathrm{rank}\,R_{\mathrm{off},k}(p)\geq r^{\sharp}$ holds for all $p\in U$ and for all $k$.
@@
The rank function is lower-semicontinuous with respect to $C^{0}$ convergence ([10]). Since $R_{\mathrm{off},k}\to R_{\mathrm{off}}$ uniformly on $M$, the inequality is preserved:
$\mathrm{rank}(R_{\mathrm{off}}(p))\geq\liminf_{k\to\infty}\mathrm{rank}(R_{\mathrm{off},k}(p))\geq r^{\sharp}\quad\text{on a dense subset of }M.$
@@
The extension of the inequality to the entire manifold $M$ follows from the fact that the set $\{p\in M\mid\mathrm{rank}\,R_{\mathrm{off}}(p)\geq r^{\sharp}\}$ is open (by lower-semicontinuity) and contains a dense subset, thus coinciding with $M$.
To extend the inequality
@@
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq r^{\sharp}$
from the dense open subset $\mathcal{U}$ to all of $M$, we use the lower-semicontinuity of the rank of a continuous family of linear maps.
The extension of the inequality $\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq r^{\sharp}$ to the entire manifold $M$ follows from the lower-semicontinuity of the rank, as established in [10]. Indeed, by Ambrose–Singer the holonomy algebra is generated by the curvature operators; since the rank of a continuous family of Lie-algebra generators is lower-semicontinuous (Warner [10]), the inequality $\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq r^{\sharp}$ extends from the dense open set to all of $M$.
@@
Finally, the lower bound depends only on $\omega]$ and on the space of $g_{1}$-parallel harmonic $2$-forms; it is therefore invariant under any deformation of $g_{2}$ and under any deformation of $g_{1}$ that preserves the space of parallel $2$-forms, hence under any product-metric deformation that leaves $\mathcal{P}_{2}$ unchanged. ∎
Remark 5.3 (Necessity of the Non-Parallel Torsion Condition).
The proof of Theorem 5.1 relies fundamentally on the non-vanishing of the linear term $R^{C}_{\mathrm{off}}\simeq\frac{1}{2}(\nabla^{\mathrm{LC}}T)_{\mathrm{off}}$, which is established via the $s\to 0$ limit of the one-parameter family of connections $\nabla^{(s)}$. This dependence on the linear term necessitates the explicit condition $\nabla^{\mathrm{LC}}T\neq 0$ for the proof to be valid.
To illustrate the necessity of this condition, we consider a case where $\nabla^{\mathrm{LC}}T=0$. If $\nabla^{\mathrm{LC}}T=0$, the curvature formula simplifies to $R^{C}=R^{\mathrm{LC}}+\frac{1}{4}Q_{T}$, where $Q_{T}$ is the term quadratic in the torsion $T$. Since $R^{\mathrm{LC}}$ is block-diagonal for a product metric, the off-diagonal curvature is solely determined by $Q_{T}^{\mathrm{off}}$. The proof of Theorem 5.1 provides no guarantee that the quadratic term $Q_{T}^{\mathrm{off}}$ alone is sufficient to generate $r^{\sharp}$ independent directions.
Consider the flat torus $M=T^{3}=T^{2}\times S^{1}$ endowed with the flat product metric $g$. Let $T$ be the constant volume form $T=dx\wedge dy\wedge dz$. This form is harmonic, non-trivial, and its mixed rank is $\text{rank}_{\mathbb{R}}([T]_{\mathrm{mixed}})=1$. Since $T$ is constant, we have $\nabla^{\mathrm{LC}}T=0$.
For this specific example, the curvature $R^{C}$ vanishes identically, leading to $\dim(\text{hol}_{p}^{\mathrm{off}}(\nabla^{C}))=0$. The lower bound is calculated as $r^{\sharp}=\operatorname{rank}_{\mathbb{R}}([T]_{\mathrm{mixed}})-\dim(\mathcal{P}_{2}\otimes H^{1}(M_{2}))=1-1=0$.
The resulting inequality $0\geq 0$ is satisfied, but the bound is trivial. This example demonstrates that when $\nabla^{\mathrm{LC}}T=0$, the topological lower bound $r^{\sharp}$ can be reduced to zero, even for a non-trivial cohomology class ($\omega]\neq 0$). The failure of the bound to provide a non-trivial constraint confirms that the contribution of the quadratic term $Q_{T}^{\mathrm{off}}$ alone is insufficient to guarantee the non-triviality of the topological lower bound, necessitating the explicit condition $\nabla^{\mathrm{LC}}T\neq 0$ for the proof of Theorem 5.1 to hold and to yield a meaningful geometric result.  [^1] The bound is recovered if we consider a non-parallel torsion $T=\omega_{h}+\varepsilon d(\sin(x)dy\wedge dz)$, which maintains the same cohomology class $\omega_{h}]$ but ensures $\nabla^{\mathrm{LC}}T\neq 0$.
@@
The invariance of the lower bound under $g_{1}$-deformations that preserve $\mathcal{P}_{2}$ is non-trivial and geometrically significant on manifolds $M_{1}$ with special holonomy (e.g., Calabi-Yau or flat tori). In these cases, the space of parallel forms $\mathcal{P}_{2}$ is non-zero, and the existence of non-homothetic deformations preserving $\mathcal{P}_{2}$ is guaranteed by known results in differential geometry (e.g., deformation theory of special holonomy metrics). This highlights the topological nature of the bound, which is stable under a non-trivial class of metric variations.
Remark 5.5 (Forced irreducibility and quantified off-diagonal holonomy).

### v9

@@ 5. Main Result and Proof
$\Psi_{p}:\mathcal{V}_{2,1}\longrightarrow\mathfrak{so}(T_{p}M),\qquad\Psi_{p}\Bigl(\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\otimes\beta_{i}\Bigr):=\frac{1}{2}\sum_{i=1}^{r_{2,1}}c_{i}\,(\nabla^{LC}T)^{\mathrm{off}}(X_{i},Y_{i},Z_{i}).$ (5.5)
The map is well‑defined and linear; its image consists precisely of the off‑diagonal curvature operators generated by the linear term $\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ arising from the (2,1)(2,1)‑component of $\omega_{h}$.
If $\sum_{i}c_{i}\alpha_{i}\otimes\beta_{i}\in\ker\Psi_{p}$, then the corresponding off‑diagonal curvature contribution vanishes:
$\frac{1}{2}\sum_{i}c_{i}(\nabla^{LC}T)^{\mathrm{off}}(X_{i},Y_{i},Z_{i})=0.$
Using the product rule $\nabla^{LC}(\alpha_{i}\wedge\beta_{i})=(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}+\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i})$ and the fact that $R^{LC}_{\mathrm{off}}=0$, this condition is equivalent to
$\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{LC}(\alpha_{i}\wedge\beta_{i})=0.$ (5.6)
Separating the bidegrees in (5.6) yields the two independent conditions
$\sum_{i=1}^{r_{2,1}}c_{i}(\nabla^{g_{1}}\alpha_{i})\wedge\beta_{i}=0,\qquad\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\wedge(\nabla^{g_{2}}\beta_{i})=0.$ (5.7)
Because the forms $\beta_{i}$ (resp. $\alpha_{i}$) are linearly independent on an open‑dense set, we obtain
$\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{g_{1}}\alpha_{i}=0,\qquad\sum_{i=1}^{r_{2,1}}c_{i}\nabla^{g_{2}}\beta_{i}=0$ (5.8)
on that set. By the unique‑continuation principle for harmonic forms (Lemma 5.1), these equalities extend globally, hence
$\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\in\mathcal{P}_{2}(M_{1}),\qquad\sum_{i=1}^{r_{2,1}}c_{i}\beta_{i}\in\mathcal{P}_{1}(M_{2}).$ (5.9)
Conversely, if (5.9) holds, then $\nabla^{LC}\bigl(\sum_{i}c_{i}\alpha_{i}\wedge\beta_{i}\bigr)=0$, which forces $\Psi_{p}\bigl(\sum_{i}c_{i}\alpha_{i}\otimes\beta_{i}\bigr)=0$. Therefore $\xi=\sum_{i}c_{i}\alpha_{i}\otimes\beta_{i}$ belongs to $\ker\Psi_{p}$ if and only if (5.9) is satisfied.
Now we identify this kernel with the intersection $\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr)$. If $\xi\in\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}\otimes\mathcal{P}_{1}\bigr)$, then $\xi$ can be written as a finite sum $\sum_{k}A_{k}\otimes B_{k}$ with $A_{k}\in\mathcal{P}_{2}$, $B_{k}\in\mathcal{P}_{1}$. For each term, $\nabla^{LC}(A_{k}\wedge B_{k})=0$ because $A_{k}$ and $B_{k}$ are parallel, hence $\Psi_{p}(A_{k}\otimes B_{k})=0$. By linearity, $\Psi_{p}(\xi)=0$, so $\xi\in\ker\Psi_{p}$.
For the converse, assume that $\xi=\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\otimes\beta_{i}\in\ker\Psi_{p}$. Equation (5.9) gives
$A:=\sum_{i=1}^{r_{2,1}}c_{i}\alpha_{i}\in\mathcal{P}_{2}(M_{1}),\qquad B:=\sum_{i=1}^{r_{2,1}}c_{i}\beta_{i}\in\mathcal{P}_{1}(M_{2}).$
Let $U=\operatorname{span}\{\alpha_{i}\}\subset\mathcal{H}^{2}(M_{1})$ and $V=\operatorname{span}\{\beta_{i}\}\subset\mathcal{H}^{1}(M_{2})$. Set $U_{\parallel}=U\cap\mathcal{P}_{2}(M_{1})$ and $V_{\parallel}=V\cap\mathcal{P}_{1}(M_{2})$. Choose any basis $\{\alpha_{1},\dots,\alpha_{s}\}$ of $U_{\parallel}=U\cap\mathcal{P}_{2}(M_{1})$ and complete it to a basis of $U$ by adding vectors $\{\alpha_{s+1},\dots,\alpha_{\dim U}\}$ whose classes are a basis of the quotient $U/U_{\parallel}$ (hence linearly independent modulo $\mathcal{P}_{2}(M_{1})$). By Lemma 5.1, at a generic point $p_{1}\in M_{1}$ the covariant derivatives $\{\nabla^{g_{1}}\alpha_{i}(p_{1})\}_{i=s+1}^{\dim U}$ are linearly independent. Since $A$ is parallel, $\nabla^{g_{1}}A\equiv 0$; evaluating at such a $p_{1}$ yields
$\sum_{i>s}c_{i}\nabla^{g_{1}}\alpha_{i}(p_{1})=0,$
and the linear independence forces $c_{i}=0$ for all $i>s$.
An identical construction applied to $M_{2}$ (choosing a basis of $V_{\parallel}$ and completing it to a basis of $V$ such that the remaining vectors form a basis of the quotient $V/V_{\parallel}$) implies $c_{i}=0$ for all $i>t$.
Hence, the only possibly non-zero coefficients are those with $i\leq\min(s,t)$. For such indices, by construction $\alpha_{i}\in\mathcal{P}_{2}(M_{1})$ and $\beta_{i}\in\mathcal{P}_{1}(M_{2})$. Therefore,
$\xi=\sum_{i=1}^{\min(s,t)}c_{i}\alpha_{i}\otimes\beta_{i}\in\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2}).$
Since $\xi\in\mathcal{V}_{2,1}$ by hypothesis, we have shown
$\ker\Psi_{p}\subseteq\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr).$
The opposite inclusion is immediate: any element of $\mathcal{V}_{2,1}$ that belongs to $\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})$ is a finite sum of simple tensors of parallel forms; consequently its covariant derivative vanishes, and hence it lies in $\ker\Psi_{p}$.
Thus we have proved
$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr).$
An identical argument applied to the $(1,2)$-component yields
$\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr).$
Now consider the induced map on the quotient
$\overline{\Psi}_{p}:\mathcal{V}_{2,1}/\ker\Psi_{p}\longrightarrow\mathfrak{so}(T_{p}M),\qquad\overline{\Psi}_{p}([\xi]):=\Psi_{p}(\xi).$
The map $\overline{\Psi}_{p}$ is well‑defined and linear. It is injective because $\overline{\Psi}_{p}([\xi])=0$ implies $\Psi_{p}(\xi)=0$, hence $\xi\in\ker\Psi_{p}$ and $\xi]=0$. Consequently, $\overline{\Psi}_{p}$ sends a basis of the quotient to a set of linearly independent off‑diagonal curvature operators, and we obtain
$\dim(\operatorname{im}\Psi_{p})=\dim\bigl(\mathcal{V}_{2,1}/\ker\Psi_{p}\bigr)=r_{2,1}-\dim\ker\Psi_{p}.$ (5.10)
An entirely analogous argument for the (1,2)(1,2)‑component gives
$\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr)$ (5.11)
and
$\dim(\operatorname{im}\widetilde{\Psi}_{p})=r_{1,2}-\dim\ker\widetilde{\Psi}_{p}.$ (5.12)
Hence, on the open‑dense set $\mathcal{U}$ where the harmonic evaluations are independent, the linear term $\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ generates at least
$\bigl(r_{2,1}-\dim\ker\Psi_{p}\bigr)+\bigl(r_{1,2}-\dim\ker\widetilde{\Psi}_{p}\bigr)=r-\dim\mathcal{K}=r^{\sharp}$
linearly independent off‑diagonal curvature operators. By the Ambrose–Singer theorem [2], these operators belong to the holonomy algebra, whence
$\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}\qquad\text{for all }p\in\mathcal{U}.$
@@
Then the off-diagonal curvature satisfies $R_{\mathrm{off}}(s)=s(\nabla^{LC}T)_{\mathrm{off}}+s^{2}Q_{\mathrm{off}}$. Differentiating at $s=0$, we see that the image of $(\nabla^{LC}T)_{\mathrm{off}}$ has dimension at least $r^{\sharp}$. The key step is to ensure that the rank of $R_{\mathrm{off}}(s)$ at $s=1$ is not reduced by the quadratic term $Q_{\mathrm{off}}$.
The off-diagonal curvature is given by $R_{\mathrm{off}}(s)=sA_{1}+s^{2}A_{2}$, where $A_{1}=\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ and $A_{2}=Q_{\mathrm{off}}$. On the open-dense set $U$ (defined in Lemma 5.1), the matrix $A_{1}$ has rank $\mathrm{rank}(A_{1})=r^{\sharp}$. The rank $r^{\sharp}$ is established by the harmonic component $\omega_{h}$ of $T$. The non-harmonic part $T-\omega_{h}=d\alpha+\delta\beta$ contributes to $A_{1}$, but since the bound is a lower bound, this contribution cannot reduce the rank below $r^{\sharp}$ on the dense set $U$ and is therefore absorbed by the overall estimate $\mathrm{rank}(A_{1})\geq r^{\sharp}$. More formally, a systematic cancellation of the off-diagonal contributions by the non-harmonic part $\eta=d\alpha+\delta\beta$ on an open set would require $\nabla\eta=-\nabla\omega_{h}$ locally. By the Unique Continuation Principle, this identity would extend globally, implying $\omega_{h}]=0$, which contradicts our hypothesis of a non-trivial mixed class. This ensures that the topological information carried by $\omega_{h}$ is rigidly encoded in the curvature.
@@
To prove that $\mathrm{rank}(R_{\mathrm{off}}(1))\geq r^{\sharp}$ for the smooth ($C^{\infty}$) metric $g$, we employ the method of analytic approximation and stability of rank (as stated in Remark 5.6).
By the density theorem of Greene–Jacobowitz [4] and the analyticity of harmonic forms (Morrey–Nirenberg [8]), we can select a sequence of real-analytic product metrics $g_{k}=g_{1,k}\oplus g_{2,k}$ converging to $g$ in the $C^{\infty}$ topology and satisfying
$\dim\mathcal{P}_{k}(M_{i},g_{k})\leq\dim\mathcal{P}_{k}(M_{i},g)\qquad(k=1,2).$
This selection is possible because the dimension function $g^{\prime}\mapsto\dim\mathcal{P}_{k}(M_{i},g^{\prime})$ is upper semicontinuous; consequently, the set $\{g^{\prime}\mid\dim\mathcal{P}_{k}(M_{i},g^{\prime})\leq\dim\mathcal{P}_{k}(M_{i},g)\}$ is open in the $C^{\infty}$ topology. Real-analytic metrics are dense in this open set by Greene–Jacobowitz. For such a sequence we have $\dim\mathcal{K}(g_{k})\leq\dim\mathcal{K}(g)$, and therefore
$r^{\sharp}(g_{k})\;:=\;\operatorname{rank}_{\mathbb{R}}\bigl([\omega]_{\mathrm{mixed}}\bigr)-\dim\mathcal{K}(g_{k})\;\geq\;r^{\sharp}(g).$
@@
For each fixed $k$, the curvature tensor $R_{\mathrm{off},k}(s)$ of the connection $\nabla^{(s)}=\nabla^{LC}(g_{k})+\frac{s}{2}T$ is real-analytic in $(s,p)$. On the dense open set $U$ furnished by Lemma 5.1 we have $\operatorname{rank}(A_{1,k})=r^{\sharp}(g_{k})$, where $A_{1,k}=\frac{1}{2}(\nabla^{LC}(g_{k})T)_{\mathrm{off}}$. Hence there exists an $r^{\sharp}(g_{k})\times r^{\sharp}(g_{k})$ minor $M_{k}(s)$ of $R_{\mathrm{off},k}(s)$ with $M_{k}(s)=s^{r^{\sharp}(g_{k})}D_{k}(s)$ and $D_{k}(0)\neq 0$. By the Identity Principle for Analytic Functions, $D_{k}(s)$ does not vanish identically and its zero set is discrete; consequently, for each $k$ there exists an open–dense subset $U_{k}\subseteq U$ such that
$\operatorname{rank}R_{\mathrm{off},k}(p)\geq r^{\sharp}(g_{k})\qquad\forall p\in U_{k}.$
Let $W:=\bigcap_{k=1}^{\infty}U_{k}$. Since $M$ is compact, it carries a complete Riemannian metric; in a complete metric space, a countable intersection of open dense subsets is dense. Thus $W$ is dense in $U$, hence dense in $M$. For every $p\in W$ and every $k$ we have
$\operatorname{rank}R_{\mathrm{off},k}(p)\geq r^{\sharp}(g_{k}).$
@@
The rank of a continuous family of linear maps is lower semicontinuous with respect to uniform convergence [9, Warner]. Since $R_{\mathrm{off},k}\to R_{\mathrm{off}}$ uniformly on $M$, for every $p\in W$ we obtain
$\operatorname{rank}R_{\mathrm{off}}(p)\;\geq\;\liminf_{k\to\infty}\operatorname{rank}R_{\mathrm{off},k}(p)\;\geq\;\liminf_{k\to\infty}r^{\sharp}(g_{k})\;\geq\;r^{\sharp}(g).$
@@
The set $\{p\in M\mid\operatorname{rank}R_{\mathrm{off}}(p)\geq r^{\sharp}(g)\}$ is open (by lower semicontinuity) and contains the dense subset $W$; therefore it is all of $M$. Thus $\operatorname{rank}R_{\mathrm{off}}(p)\geq r^{\sharp}(g)$ for every $p\in M$.
Finally, by Definition 3.2 the off‑diagonal holonomy algebra $\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})$ is the Lie subalgebra generated by the off‑diagonal curvature operators and their covariant derivatives at $p$. In particular, it contains the linear span of the off‑diagonal curvature operators $R_{\mathrm{off}}(p)$. Therefore,
@@
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;\operatorname{rank}R_{\mathrm{off}}(p)\;\geq\;r^{\sharp}(g)\qquad\forall p\in M,$
@@
which is the desired bound. The lower bound depends only on $\omega]$ and on the spaces of $g_{1}$‑parallel harmonic 22‑forms and 11‑forms, and of $g_{2}$‑parallel harmonic 11‑forms and 22‑forms; it is therefore invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these parallel‑form strata, hence under any product‑metric deformation that leaves $\mathcal{P}_{k}(M_{i})$ unchanged for $k=1,2$. ∎
Remark 5.3 (Necessity of the Non‑Parallel Torsion Condition).
Theorem 5.2 yields a non‑trivial lower bound only when the linear term
$\tfrac{1}{2}(\nabla^{\mathrm{LC}}T)_{\mathrm{off}}$
is non‑zero. If instead $\nabla^{\mathrm{LC}}T=0$, the curvature reduces to
$R^{C}=R^{\mathrm{LC}}+\tfrac{1}{4}Q_{T},$
and there is no guarantee that the purely quadratic off‑diagonal part $(Q_{T})_{\mathrm{off}}$ generates $r^{\sharp}$ independent directions.
Consider for instance the flat torus $M=T^{2}\times S^{1}$ with the constant volume form $T=dx\wedge dy\wedge dz$, which is harmonic and satisfies
$\operatorname{rank}_{\mathbb{R}}([T]_{\mathrm{mixed}})=1.$
Since $T$ is parallel, $\nabla^{\mathrm{LC}}T=0$; hence every simple tensor in any minimal harmonic decomposition is completely parallel. In the notation of Theorem 5.2 we have
$\mathcal{V}_{2,1}=\operatorname{span}\{dx\wedge dy\otimes dz\},\qquad\mathcal{P}_{2}(T^{2})=\operatorname{span}\{dx\wedge dy\},\qquad\mathcal{P}_{1}(S^{1})=\operatorname{span}\{dz\},$
so that
$\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(T^{2})\otimes\mathcal{P}_{1}(S^{1})\bigr)=\mathcal{V}_{2,1},\qquad\ker\widetilde{\Psi}_{p}=0.$
Thus $\dim\mathcal{K}=1$ and consequently
$r^{\sharp}=\operatorname{rank}_{\mathbb{R}}([T]_{\mathrm{mixed}})-\dim\mathcal{K}=1-1=0.$
A direct computation gives $R^{C}=0$, hence
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})=0,\qquad r^{\sharp}=0.$
The bound $0\geq 0$ is saturated but trivial, illustrating that without the non‑parallel torsion condition the theorem cannot force irreducibility.
@@
The invariance of the lower bound $r^{\sharp}$ is a non‑trivial and geometrically significant property. The bound depends only on two pieces of data:
  • the mixed class $\omega]_{\mathrm{mixed}}$,
  • the parallel‑form strata $\mathcal{P}_{k}(M_{i})$ of the factor metrics.
Explicitly,
$\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p},\qquad\ker\Psi_{p}=\mathcal{V}_{2,1}\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),\qquad\ker\widetilde{\Psi}_{p}=\mathcal{V}_{1,2}\cap\bigl(\mathcal{P}_{1}(M_{1})\otimes\mathcal{P}_{2}(M_{2})\bigr),$
where $\mathcal{V}_{2,1}\subset\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2})$ (resp. $\mathcal{V}_{1,2}$) denotes the linear span of the harmonic representatives of the $(2,1)$- (resp. $(1,2)$-) Künneth component of $\omega]$.
Consequently, $r^{\sharp}$ is invariant under any deformation of the product metric $g=g_{1}\oplus g_{2}$ that preserves the parallel‑form strata $\mathcal{P}_{k}(M_{i})$ ($k=1,2$). Such deformations include, for instance, any variation of $g_{i}$ inside the connected component of the moduli space of metrics with a fixed set of parallel forms.
This stability is especially relevant when at least one factor $M_{i}$ possesses special holonomy (e.g., Calabi–Yau manifolds or flat tori), where the spaces $\mathcal{P}^{k}(M_{i})$ are non‑trivial and the metric $g_{i}$ admits large families of deformations that leave $\mathcal{P}^{k}(M_{i})$ unchanged.
Thus, $r^{\sharp}$ quantifies the part of the topological “mixing strength” of $\omega]$ that cannot be “absorbed” by metric deformations respecting the parallel‑form strata; it captures a genuinely topological obstruction to reducible holonomy.
Remark 5.5 (Forced irreducibility and quantified off‑diagonal holonomy).